An elevator weighs 2500 𝑙𝑏; a cable, which can sustain a maximum tension of 6000 𝑙𝑏, supports it. If the elevator is going down with a velocity of 9 𝑓𝑑/𝑠, find the minimum safe distance it can travel before coming to a stop.


1
Expert's answer
2021-10-05T10:08:00-0400

Find the maximum acceleration that will create a 6000-lb-force on stop:


"ma=-mg+T,\\
a=\frac Tm-g,\\
d=\frac{v^2}{2a}=\frac{v^2m}{2T-mg}=0.9\text{ ft}."